Beef 4 capital goods and 2l Complete the following (approximate) possibilities for Emilon: a) 130 rice and to PP2? b) rice and 40 beef Which of the following pssibilities is Emilon capable of producing? c) 100 rice and 40 beef d) 70 rice and 35 beef 3. (LO 7) Table 1.14 shows the production possibilities for the country of Emilon. TABLE 1.14 30 Rice 0 50 90 120 140 Beef 50 4535 20

THE ECONOMIC PROBLEM CHAPTER 1 a What is the total cost of producing 90 rice? er(rice/beef) b What is the total cost of going from possibility C to possibility D? c What is the approximate per-unit cost of going from possibility C to possibility D d What is the total cost of going from possibility D to possibility C? Answer: Answer:(rice/beef Answer(rice/beef) Answer (rice/beef) Intermediate (Problems 4-7)
Answer #1

Question 3

When rice production is increased from 0 to 90, the production of beef is reduced from 50 to 35 so the cost of producing 90 rice is 15 beef or 15/90 = 0.167 beef per unit of rice.

From C to D, there is a reduction in the production of beef by 15. Hence the total cost is 15 beef

Per unit cost is 15/(120 - 90) = 0.5 beef per unit of rice.

From D to C there is a reduction in the production of rice by 30. Hence the total cost is 30 rice or 30/15 = 2 rice per beef.